1.) Given P(A) = 0.57, P(B) = 0.20, P(C) = 0.34 and that events A, B, and C are independent, what is P(A, B, and C).

Answer in decimal form. Round to 3 decimal places as needed.
Answer #1

Given that A, B and C are independent. Hence,

P(A, B and C) = P(A) * P(B) * P(C) = 0.57 * 0.20 * 0.34 = 0.039